OpenStudy (anonymous):

PLease Help!! Find the complex solutions of the quadratic equation: 3(x^2)+2x+18=0

OpenStudy (campbell_st):

ok... so you will need the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] now in your question a = 3, b = 2 and c = 18 you'll need to substitute them and then evaluate to get the complex solutions... it may be possible to simplify the solutions.

OpenStudy (anonymous):

I tried that and we supposed to use an imaginary number i . I ve entered my answer 3 times but I still get it wrong, it is an online website for homework problems.

OpenStudy (campbell_st):

yes... you'll get to that... what did you get when you substituted.... and evaluated...

OpenStudy (anonymous):

i got for the final answer: (-1+2(sqrt53)i)/3 and (-1-2(sqrt53)i)/3

OpenStudy (campbell_st):

so its \[x = \frac{ -2 \pm \sqrt{-212}}{6}\] now \[i^2 = -1\] so you will have \[x = \frac{-2 \pm \sqrt{212i^2}}{6}\]

OpenStudy (campbell_st):

so this will simplify to \[x = \frac{-1 \pm i \sqrt{53}}{3}\] hope this makes sense

OpenStudy (anonymous):

Thanks, the website finally accepted the answer.

OpenStudy (campbell_st):

glad to help